### v1

@@ Abstract.
The field of collective intelligence studies how teams can achieve better results than any of the team members alone. The special case of human-machine teams carries unique challenges in this regard. For example, human teams often achieve synergy by communicating to discover their relative advantages, which is not an option if the team partner is an unexplainable deep neural network. Between 2005-2008 a set of ”freestyle” chess tournaments were held, in which human-machine teams known as ”centaurs”, outperformed the best humans and best machines alone. Centaur players reported that they identified relative advantages between themselves and their chess program, even though the program was superhuman. Inspired by this and leveraging recent open-source models, we study human-machine like teams in chess. A human behavioral clone (”Maia”) and a pure self-play RL-trained chess engine (”Leela”) were composed into a team using a Mixture of Experts (MoE) architecture. By directing our research question at the selection mechanism of the MoE, we could isolate the issue of extracting relative advantages without knowledge sharing. We show that in principle, there is high potential for synergy between human and machine in a complex sequential decision environment such as chess. Furthermore, we show that an expert can identify only a small part of these relative advantages, and that the contribution of its subject matter expertise in doing so saturates quickly. This is probably due to the ”curse of knowledge” phenomenon. We also train a network to recognize relative advantages using reinforcement learning, without chess expertise, and it outdoes the expert. Our experiments are repeated in asymmetric teams, in which identifying relative advantages is more challenging. Our findings contribute to the study of collective intelligence and human-centric AI.
@@ 1. Introduction
Collective intelligence studies how teams can outperform any one of their members. The literature has shown that in human teams, a moderate amount of diversity and open communication among members, are correlated with team synergy (Malone,, 2019). However, these conditions have limited applicability in human-machine teams, which have sharp cognitive diversity and which are communication impaired, especially in the case of unexplainable deep neural networks. We focus on the task of discovering relative advantages between team members in sequential decision making, and ask how it can be accomplished under these conditions. In particular, we are interested in the role of domain knowledge in making this judgement. For example, in management, there is a debate about whether it is beneficial for business managers to be subject matter experts to successfully manage their teams (Phillips, M.,, 2011).
To answer these questions, we take inspiration from a set of eight ”freestyle” chess tournaments held between 2005-2008 (Chess Base,, 2007), in which participants could play either as a human, a chess program or as a ”centaur”. In a centaur team, a human and machine work together analyzing the board positions, but the human makes the ultimate decisions. The highest rank in these tournaments was consistently taken by a centaur team, despite the fact that the chess programs at the time were already superhuman. Centaur players have reported that knowing the relative advantages of their machine counterparts is an important element in their success (Chess Base,, 2007).
This claim has an unintuitive implication: that the cognitive function of identifying relative advantages is not dependent on having superior domain knowledge to the team members. (Krakowski et al.,, 2022) have gone so far as to argue that in asymmetric human-machine teams, subject matter skills of the human aren’t important at all, and that instead complementary human skills such as configuring the machine are the true sources of synergy.
To study this question, we performed computational experiments that leverage the recent availability of open-source chess models trained under different regimes: a human-like behavioral clone trained with supervised learning from human games only, and a pure self-play RL agent with zero exposure to human data. The behavioral clone and the pure self-play RL agent are composed into a Mixture of Experts (MoE) model (Yuksel et al.,, 2012), which allows us to simulate human-machine teaming at scale, albeit in a simplified lab setting. This also allows us to isolate the function of finding relative advantages, in the selection mechanism of the MoE.
@@ 2.1. Team Players
We deploy two types of chess models to play as a team. One is Leela-Chess-Zero (”Leela”), an open-source version of Alpha-Zero (Silver et al.,, 2017), which is a pure self-play reinforcement learning algorithm. Due to its lack of exposure to human data or knowledge, it has been noted for its non-human or ”alien” style of play (Manella and Zohar,, 2021). The other is Maia-Chess (”Maia”), an open-source family of models designed to be human-like, trained exclusively from human games (McIlroy-Young et al.,, 2020). Maia and Leela play as a team against a popular chess engine known as Stockfish, which has a heuristics-based evaluation function, parameterized by a shallow neural network.
@@ 2.2. Team Model
This setup can be formally described as follows. In the RL literature (Sutton and Barto,, 2020), Markov Decision Processes (MDP) are considered, consisting of states S, actions A, a transition function T, a reward function R, and a discount factor $\gamma$ giving the tuple (S, A, T, R, $\gamma$). A policy $\pi$ chooses actions given states. Expected accumulated reward of a policy, starting from a state-action pair, is known as the Q function of the policy.
@@
In this model, the original human centaur player is conceived as having two distinct functions: forming a substantive opinion about the best action (as one of the base policies) and estimating relative advantages (as the manager). In other domains, one could also envisage an automated manager that gives a relative advantage score to team members in each state.
@@ 2.3. RL Trained Manager
where s is the position being evaluated, $a_{M}$ and $a_{L}$ are the move recommendations in s and $\text{RLManager}_{i}$ is the RL trained manager at iteration i. $Q_{\hat{\pi}}$ represents the Q value for the team given the action recommendations $a_{M}$ and $a_{L}$, estimated empirically by the outcome of the game rollouts from s. The resulting dataset was then used to train $\text{RLManager}_{i+1}$. At inference, the RL manager classifies the player to listen to, without performing game rollouts.
@@
The RL manager was parameterized by a transformer architecture (illustraion in figure 2). As input, the model receives a representation of the board using 64 tokens, one for each square. The tokens denote the pieces that occupy each square, or if it is empty, plus positional encoding. Additional tokens signify the color being played, castling rights and whether the king is in check. Following BERT (Devlin et al.,, 2019), there is also an auxiliary ”CLS” token. The CLS token in the last layer is fed to a fully-connected classifier network which outputs a binary indicator for Maia or Leela.
Note that the RL manager receives only the state as input, without move recommendations. If it were exposed to the move recommendations, it may learn to play in its own right, which we wish to control for. A network trained in this manner represents a manager that has learned relative advantages between team members using statistical methods only, without independent skill in the task.
@@
Chess engines are developed using games starting from diverse opening positions, to increase coverage of the game space. We collected over 28000 chess opening positions for training games.
@@ 2.4. Expert Manager
To examine the role of domain knowledge in finding relative advantages, we use an external strong subject matter ”expert” chess engine as the manager. The expert has access to the recommended moves of Maia and Leela, and it scores out each of them using its own evaluation function. Formally, if we define a subject matter expert policy $\pi_{SME}$, we can define an expert-controlled team as:
@@
Note also that in a non-sequential decision task, such an expert is theoretically the best possible manager. In a sequential decision task this is less clear. This is because the assumption when evaluating $Q_{\pi_{SME}}$ is that the subject matter expert $\pi_{SME}$ continues alone after the chosen move, which is not what happens in the team. For the expert manager, we used a strong version of Stockfish (v. 14), experimenting with different search depths.
@@ 2.5. Oracle Manager
To compare the performance of these managers to the maximal potential synergy in a given Maia-Leela pair, we can posit an additional ”oracle” manager. It has online access to the team player recommendations and the adversary responses, simulates all trajectories in the game tree from each decision point and finds the optimal decision using backward-induction. This would be computationally demanding due to the burden of opening out full game trees and obtaining each player’s move recommendations at every node. Alternatively, we conceptualize an approximate oracle that simulates each of the recommended actions, and thereafter continues just these two branches using the better of the two team players, against the fully modelled adversary. The oracle then returns to the original decision point and chooses the player whose recommended move led to the better outcome in the simulation. Since this only approximates the full theoretical oracle, it will not produce the optimal team performance, but it would provide a lower bound for that optimum. The approximate oracle, while more computationally feasible, is still not an admissible manager online during games, since it has full access to the adversary policy. However, comparing the other managers to this approximate oracle manager, gives some indication of the remaining headroom for synergy.
@@ 2.6. Evaluation
The baseline for synergy is the WDL of the best team member playing by itself against the adversary ($max(WDL_{\pi_{M}},WDL_{\pi_{L}})$).
@@ 3.1. Results for a Symmetric Team
First we examine a symmetric team, i.e. where the Leela network is almost the same strengh as the Maia network, playing against a Stockfish (v. 11) with search depth=1. Results are displayed in figure 3.
@@
To more closely examine the role of subject matter expertise, the manager was then deployed at increasing levels of search depth (3, 5 and 15). At depth 15, if the expert was playing against the adversary itself without being constrained to Maia or Leela recommendations, it would have WDL score of 0.942. All versions of the subject matter expert achieve significant team synergy, but increasing its strength does not yield much improvement (WDL at peak was 0.501). This seems to indicate that the marginal contribution of domain knowledge to finding relative advantages saturates quickly. The limitations of the expert as manager can be viewed as a demonstration of the ”curse of knowledge” (Camerer et al.,, 1989). Intuitively, if the manager is too smart, it may not properly account for the future behavior of the inferior team players after the current move.
@@
The oracle far outperforms other managers, achieving WDL of 0.799, indicating vast headroom for synergy not found by the other managers.
@@ 3.2. Results for Asymmetric Teams
We repeated the above experiment for three different levels of asymmetry within the Maia-Leela team. The Maia network was held fixed, while successively stronger Leela networks were chosen to team with it. The Stockfish adversary (v. 14) was also successively improved by setting it’s search depth to levels 4, 7 and 10 respsectively. Leela and Stockfish were callibrated to be approximately on par with each other in each experiment, that is Leela has a WDL consistently close to 0.5. The subject matter expert used in all teams was Stockfish (v. 14) at search depth=15.
@@
Figure 4. WDL for asymmetric teams. Synergy is above the WDL of Leela alone. Error bars represent SEM.
@@ 3.3. Distribution of Team Member Choices
Results are show in figure 5. The figure shows that the oracle tends to extremes, especially in asymmetric teams. At first, this may seem to be due to the bias introduced by simulating the continuation after each move recommendation using one fixed team member. However, when considering the high performance of the oracle, it must have identified the most consequential contributions by the inferior player. If we infer from this that there is only a small kernel of high-impact decisions, it becomes apparent why finding them is hard for the other managers.
Figure 5. The oracle chooses the inferior player more sparingly than the other managers, indicating the existence of a small kernel of high-impact decisions.
@@ 4. Explaining the RL Manager
Recall that we trained the RL manager without exposing it to any rules of chess or to the recommended moves of the team players. It was only fed static positions and binary Maia/Leela labels. We compared it to the expert manager since we wanted to know if domain knowledge is decisive in the task of finding relative advantages. However, it is possible that the RL manager developed chess knowledge implicitly, as well as knowledge of the team members’ policies, and uses these to identify the best recommendation. That is, it might have learned to become a quasi subject matter expert.
@@ 4.1. Is domain knowledge necessary to identify relative advantage?
First, we ask whether the RL manager uses non-trivial understanding of the domain to make its decisions. We explore this using the attention scores in the transformer network. Attention scores are used in transformer models to weight input tokens (in our case squares on the chess board) during processing, which is sometimes leveraged for explainability (Vaswani et al.,, 2017; Wiegreffe and Pinter,, 2019).
To see whether the RL manager learns something about the rules of chess, we checked if the network gives more attention to pieces than to empty squares. Over 8000 chess positions were fed to the network. Attention scores of the CLS token towards tokens representing squares on the board were recorded (averaged over layers and attention heads). For each position the mean attention to pieces and to empty squares were taken. For control, we compared to the same attentions given by an untrained network. Since pieces are correlated with regions of the board, we used an additional control - the trained network attentions over shuffled chess positions.
Results are shown in figure 6(c). In the case of the untrained network, attentions are near uniform. With the shuffled data, the trained network gives more attention on average to pieces than empty squares. With the correctly ordered data, the difference in attention to pieces and to empty squares is much more substantial, to the point that the two distributions hardly overlap.
@@
Results are shown in figure 6(d). As before, the untrained network gives near-uniform attention to all squares. With the shuffled data, there is no preference for attacked pieces over non-attacked. But over the ordered data the trained network clearly gives more attention to attacked pieces than not. We measure effect size using non-parametric $A_{w}$ which is a robust test for group differences in numeric variables (Li et al.,, 2021). It gives a high effect size of 0.81 (1 is highest).
Figure 6. a) Illustration of the attentions extracted from the model; b) Attention heatmap overlays (lighter colors denote higher attention); c) Comparison of attentions to pieces vs empty squares; d) Comparison of attentions to attacked pieces vs not-attacked. Whiskers in the box plots represent farthest data point lying within 1.5x the inter-quartile range from the box.
The network focuses more on pieces than empty squares, and more on threatened pieces than non-threatened. We don’t know how deep the domain knowledge goes, but at least, the model locates the pieces and observes their interrelationships under the rules of chess. This indicates that the manager obtained some domain knowledge while learning Maia and Leela relative advantages.

### v2

@@ Abstract.
The field of collective intelligence studies how teams can achieve better results than any of the team members alone. The special case of human-machine teams carries unique challenges in this regard. For example, human teams often achieve synergy by communicating to discover their relative advantages, which is not an option if the team partner is an unexplainable deep neural network. Between 2005-2008 a set of ”freestyle” chess tournaments were held, in which human-machine teams known as ”centaurs”, outperformed the best humans and best machines alone. Centaur players reported that they identified relative advantages between themselves and their chess program, even though the program was superhuman. Inspired by this and leveraging recent open-source models, we study simulated human-machine teams in chess. A human behavioral clone (”Maia”) and a pure self-play RL-trained chess engine (”Leela”) were composed into a team using a Mixture of Experts (MoE) architecture. By directing our research question at the selection mechanism of the MoE, we could isolate the issue of extracting relative advantages without knowledge sharing. We show that in principle, there is high potential for synergy between human and machine in a complex sequential decision environment such as chess. Furthermore, we show that an expert can identify only a small part of these relative advantages, and that the contribution of its subject matter expertise in doing so saturates quickly. This is probably due to the ”curse of knowledge” phenomenon. We also train a network to recognize relative advantages using reinforcement learning, without chess expertise, and it outdoes the expert. Our experiments are repeated in asymmetric teams, in which identifying relative advantages is more challenging. Our findings contribute to the study of collective intelligence and human-centric AI.
@@ 1. Introduction
Collective intelligence studies how teams can outperform any one of their members. The literature shows that in human teams, a moderate amount of diversity and open communication among members, are correlated with team synergy Malone (2019). However, these conditions have limited applicability in human-machine teams, which have sharp cognitive diversity and which are communication-impaired, especially in the case of unexplainable deep neural networks. We study the task of discovering relative advantages between team members in sequential decision making, and ask how it can be accomplished under these conditions. In particular, we are interested in the role of domain expertise in making this judgment.
To answer these questions, we take inspiration from a set of eight ”freestyle” chess tournaments held between 2005-2008 Nickel (2008), in which participants could play either as a human, a chess program or as a ”centaur”. In a centaur team, a human and machine work together analyzing the board positions, but the human makes the ultimate decisions. The highest rank in these tournaments was consistently taken by a centaur team, despite the fact that the chess programs at the time were already superhuman. Centaur players have reported that knowing the relative advantages of their machine counterparts is an important element in their success Nickel (2008).
This claim has an unintuitive implication: that the cognitive function of identifying relative advantages is not dependent on having superior domain expertise to the team members. Krakowski et al. (2022) have gone so far as to argue that in asymmetric human-machine teams, subject matter skills of the human aren’t important at all, and that instead complementary human skills such as configuring the machine are the true sources of synergy.
To study this question, we performed computational experiments that leverage the recent availability of open-source chess models trained under different regimes: a human-like behavioral clone trained with supervised learning from human games only, and a pure self-play RL agent with zero exposure to human data. The behavioral clone and the pure self-play RL agent are composed into a Mixture of Experts (MoE) model Yuksel et al. (2012), which allows us to simulate human-machine teaming at scale, albeit in a simplified lab setting. This also allows us to isolate the function of finding relative advantages, in the selection mechanism of the MoE.
@@ 2.1. Team Players
We deploy two types of chess models to play as a team. One is Leela-Chess-Zero (”Leela”), an open-source version of Alpha-Zero Silver et al. (2017), which is a pure self-play reinforcement learning algorithm. Due to its lack of exposure to human data or knowledge, it has been noted for its non-human or ”alien” style of play Manella and Zohar (2021). The other is Maia-Chess (”Maia”), an open-source family of models designed to be human-like, trained exclusively from human games McIlroy-Young et al. (2020). Maia and Leela play as a team against a popular chess engine known as Stockfish, which has a heuristics-based evaluation function, parameterized by a shallow neural network.
@@ 2.2. Team Model
This setup can be formally described as a Markov Decision Process (MDP)Sutton and Barto (2020), consisting of states S, actions A, a transition function T, a reward function R, and a discount factor $\gamma$ giving the tuple (S, A, T, R, $\gamma$). A policy $\pi$ chooses actions given states. Expected accumulated reward of a policy, starting from a state-action pair, is known as the Q function of the policy.
@@
In this model, the original human centaur player is conceived as having two distinct functions: forming a substantive opinion about the best action (as one of the base policies) and estimating relative advantages (as the manager). In other settings, one could also envisage an automated manager that gives a relative advantage score to team members in each state.
@@ 2.3. RL Trained Manager
where s is the position being evaluated, $a_{M}$ and $a_{L}$ are the move recommendations in s and $\text{RLManager}_{i}$ is the RL trained manager at iteration i. $Q_{\hat{\pi}}$ represents the Q value for the team given the action recommendations $a_{M}$ and $a_{L}$, estimated empirically by the outcome of the game rollouts from s. The resulting dataset was then used to train $\text{RLManager}_{i+1}$.
@@
The RL manager was parameterized by a transformer architecture (illustration in figure 2). As input, the model receives a representation of the board using 64 tokens, one for each square. The tokens denote the pieces that occupy each square, or if it is empty, plus positional encoding. Additional tokens signify the color being played, castling rights and whether the king is in check. Following BERT Devlin et al. (2019), there is also an auxiliary ”CLS” token. The CLS token in the last layer is fed to a fully-connected classifier network which outputs a binary indicator for Maia or Leela.
Note that the RL manager receives only the state as input, without move recommendations. This is to train a manager focused on identifying the relative advantages of the team members, rather than judging the merits of the particular recommendations which would require deep domain expertise.
@@
[…]
Transformer model architecture for chess manager
Chess engines are often developed and evaluated on games starting from diverse opening positions, to increase coverage of the game space. We collected over 28000 chess opening positions for training games.
@@ 2.4. Expert Manager
To examine the role of domain expertise in producing team synergy, we use an external strong ”subject matter expert” chess engine as the manager. The expert has access to the recommended moves of Maia and Leela, and it scores out each of them using its own evaluation function. Formally, if we define a subject matter expert policy $\pi_{SME}$, we can define an expert-controlled team as:
@@
Note also that in a non-sequential decision task, such an expert is the best possible manager. In a sequential decision task this is less clear. This is because the assumption when evaluating $Q_{\pi_{SME}}$ is that the subject matter expert $\pi_{SME}$ continues alone after the chosen move, which is not what happens in the team. For the expert manager, we used a strong version of Stockfish (v. 14), experimenting with different search depths.
@@ 2.5. Oracle Manager
To compare the performance of these managers to the maximal potential synergy in a given Maia-Leela pair, we can posit an additional ”oracle” manager, that has online access to the team player and adversary policies. At each decision, it opens out full game trees, obtains each player’s future move recommendations and adversary responses, and backward induces the best current recommendation. This would not be an admissible manager during games, since it has full access to the adversary policy for the simulations. But it would theoretically represent the highest attainable synergy within the team, allowing us to see how much the other managers fall short.
Practically, such an oracle is computationally demanding. As an alternative, we conceptualize an approximate oracle that simulates each of the recommended moves, and thereafter continues these two simulated branches with just one trajectory each, generated by continuing with the better of the two team players against the fully modeled adversary. This approximate oracle then returns to the original decision point and chooses the player whose recommended move led to the better outcome in the simulation. Since this only approximates the full theoretical oracle, it will not necessarily produce the optimal team performance, but it would provide a lower bound for that optimum.
@@ 2.6. Evaluation
The baseline for synergy is the WDL of the best team member playing by itself against the adversary: max($\text{WDL}_{\pi_{M}},\text{WDL}_{\pi_{L}}$).
@@ 3.1. Results for a Symmetric Team
First we examine a symmetric team, i.e. where the Leela network is almost the same strength as the Maia network, playing against Stockfish (v. 11) with search depth=1. Results are seen in figure 3.
@@
[…]
Results for approximately symmetric centaur team using different managers
@@
To more closely examine the role of subject matter expertise, the manager was then deployed at increasing levels of search depth (3, 5 and 15). At depth 15, if the expert was playing against the adversary itself without being constrained to Maia or Leela recommendations, it would have WDL score of 0.942. All versions of the subject matter expert achieve significant team synergy, but increasing the manager’s strength does not yield much improvement (team WDL at peak was 0.501). This seems to indicate that the marginal contribution of domain knowledge to finding relative advantages saturates quickly. The limitations of the expert as manager can be viewed as a demonstration of the ”curse of knowledge” Camerer et al. (1989). Intuitively, if the manager is too smart, it may not properly account for the future behavior of the inferior team players after the current move.
@@
The oracle achieved WDL of 0.799 far outperforming the other managers, indicating vast headroom for synergy.
@@ 3.2. Results for Asymmetric Teams
We repeated the above experiment for three different levels of asymmetry within the Maia-Leela team. The Maia network was held fixed, while successively stronger Leela networks were chosen to team with it. The Stockfish adversary (v. 14) was also successively improved by setting it’s search depth to levels 4, 7 and 10 respectively. Leela and Stockfish were calibrated to be approximately on par with each other in each experiment, that is Leela has a WDL consistently close to 0.5. The subject matter expert used in all teams was Stockfish (v. 14) at search depth=15.
@@
Figure 4. WDL for asymmetric teams. Synergy is above the WDL of Leela alone (dashed line). Error bars represent SEM.
[…]
Results for asymmetric teams using different managers
@@ 3.3. Distribution of Team Member Choices
Results are show in figure 5. Focusing on the oracle, it tends to choose the inferior player quite sparingly. This may be due to the bias introduced by simulating the continuation after each move recommendation using the superior team member. Nonetheless, when considering the high performance of the oracle, it seems to have identified the most consequential contributions by the inferior player. We can infer that there is a small kernel of high-impact decisions, which makes it apparent why finding them is hard for the other managers.
Figure 5. Distribution of choices between Maia and Leela. The oracle chooses the inferior player (Leela in the symmetric team and Maia in the others) more sparingly than the other managers.
[…]
Distribution of choices between Maia and Leela by different managers in different teams
@@ 4. Explaining the RL Manager
Recall that we trained the RL manager without exposing it to any rules of chess or to the recommended moves of the team players. It was only fed static positions and binary Maia/Leela labels. We compared it to the expert manager since we wanted to know if domain knowledge is decisive in the task of finding relative advantages. However, it is possible that the RL manager developed chess knowledge implicitly, as well as knowledge of the team members’ policies, and uses these to identify the best recommendation. That is, it might have learned to become a subject matter expert.
@@ 4.1. Is domain knowledge necessary to identify relative advantage?
First, we ask whether the RL manager uses non-trivial understanding of the domain to make its decisions. We explore this using the attention scores in the transformer network. Attention scores are used in transformer models to weight input tokens (in our case squares on the chess board) during processing, which is sometimes leveraged for explainability Vaswani et al. (2017); Wiegreffe and Pinter (2019).
To see whether the RL manager learns something about the rules of chess, we checked if the network gives more attention to pieces than to empty squares. Over 8000 chess positions were fed to the network. Attention scores of the CLS token towards tokens representing squares on the board were recorded (averaged over layers and attention heads). For each position the mean attention to pieces and to empty squares were taken. For control, we compared to the same attentions given by an untrained network. In the untrained network, attentions were near uniform. In the trained network, there was a clear difference in attention to pieces and to empty squares, to the point that the two distributions hardly overlap.
Since pieces are correlated with regions of the board, we used an additional control - the trained network attentions over shuffled chess positions. In this control, the network gives more attention on average to pieces than empty squares, but not as pronounced as in the properly ordered data. Results are shown in figure 6(c).
@@
Results are shown in figure 6(d). As before, the untrained network gives near-uniform attention to all squares. With the shuffled data, there is no preference for attacked pieces over non-attacked. But over the ordered data the trained network clearly gives more attention to attacked pieces. We measure effect size using non-parametric $A_{w}$ which is a robust test for group differences in numeric variables Li et al. (2021). It gives a high effect size of 0.81 (1 is highest).
Figure 6. Investigation of the RL-manager network. (a) Illustration of attentions extracted from the model; (b) Example attention heatmap overlay (lighter colors denote higher attention); (c) Attention to pieces vs empty squares; (d) Attention to attacked pieces vs not-attacked. Whiskers in the box plots represent farthest data point lying within 1.5x the inter-quartile range from the box.
[…]
Investigation of attention scores under rules of chess
The network focuses more on pieces than empty squares, and more on threatened pieces than non-threatened. This indicates that the manager obtained some domain knowledge while learning Maia and Leela relative advantages. We don’t know how deep the domain knowledge goes, but at least, the model locates the pieces and observes their interrelationships under the rules of chess.